C= 200 + 0.5Y I= 800 G= 600 X= 400 M = 300. Where: G is government expenditure, X is exports, M is imports, I is planned investment and C is household consumption.

what is the value of balance of trade?
1
Expert's answer
2020-03-16T12:10:21-0400

balance of trade"=X-M;"

balance of trade "=400-300=100;"